# Safe Navigation in Uncertain Crowded Environments Using Risk Adaptive CVaR Barrier Functions

### v2

@@ Abstract
Robot navigation in dynamic, crowded environments poses a significant challenge due to the inherent uncertainties in the obstacle model. In this work, we propose a risk-adaptive approach based on the Conditional Value-at-Risk Barrier Function (CVaR-BF), where the risk level is automatically adjusted to accept the minimum necessary risk, achieving a good performance in terms of safety and optimization feasibility under uncertainty. Additionally, we introduce a dynamic zone-based barrier function which characterizes the collision likelihood by evaluating the relative state between the robot and the obstacle. By integrating risk adaptation with this new function, our approach adaptively expands the safety margin, enabling the robot to proactively avoid obstacles in highly dynamic environments. Comparisons and ablation studies demonstrate that our method outperforms existing social navigation approaches, and validate the effectiveness of our proposed framework.
@@ III-A Discrete-Time Control Barrier Functions
where $\mathbf{x}_{k}=[\mathbf{p}_{k},\mathbf{v}_{k}]^{T}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state at time step $k\in\mathbb{Z}^{+}$, with $\mathbf{p}_{k}\in\mathbb{R}^{d}$ and $\mathbf{v}_{k}\in\mathbb{R}^{d}$ denoting the position and velocity of the robot w.r.t. a global frame (with $n=2d$ for a $d$-dimensional space), respectively, and $\mathbf{u}_{k}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, with $\mathcal{U}$ being the set of admissible controls for system (1). The function $f:\mathcal{X}\times\mathcal{U}\to\mathbb{R}^{n}$ is assumed to be locally Lipschitz continuous with respect to $\mathbf{x}_{k}$ and $\mathbf{u}_{k}$.
@@ III-B Probabilistic Constraints
Due to the inherent uncertainty in the dynamic obstacle scenarios, the safe states of the robot cannot be computed deterministically. Therefore, a probabilistic formulation of safety is required. Let $\mathbf{x}^{o}_{k}\in\mathcal{O}\subset\mathbb{R}^{n}$ denote the state of the obstacle at time $k$, where $\mathbf{x}^{o}_{k}=[\mathbf{p}^{o}_{k},\mathbf{v}^{o}_{k}]^{T},$ with $\mathbf{p}^{o}_{k}\in\mathbb{R}^{d}$ and $\mathbf{v}^{o}_{k}\in\mathbb{R}^{d}$ representing the position and velocity, respectively. We define the deterministic prediction model $g:\mathbb{R}^{n}\to\mathbb{R}^{n},$ which maps the current obstacle state $\mathbf{x}^{o}_{k}$ to a prediction of its next state (in the absence of noises). Therefore, the predicted next state with noises for given $\mathbf{x}^{o}_{k}$ can be represented as
$\hat{\mathbf{x}}^{o}_{k+1}=g(\mathbf{x}^{o}_{k})+\mathbf{w}_{k},$ (5)
where the noise $\mathbf{w}_{k}\in\mathbb{R}^{n}$ is a random vector representing the uncertainty in the obstacle dynamics at time $k$,
Let us define the safe set given the obstacle states as
$\mathcal{S}=\{\mathbf{x}_{k}\in\mathcal{X}:h(\mathbf{x}_{k},\mathbf{x}^{o}_{k})\geq 0\}.$ (6)
$h:\mathcal{X}\times\mathcal{O}\to\mathbb{R}$ is the CBF that depends on both the system state $\mathbf{x}_{k}$ and the obstacle state $\mathbf{x}^{o}_{k}$, where we assume perfect measurements of the states at the current time step $k$. However, the CBF at the next time step $k+1$ considering the predicted state, which is denoted as $h_{k+1}\coloneqq h\big{(}\mathbf{x}_{k+1},\hat{\mathbf{x}}^{o}_{k+1}\big{)},$ becomes a random variable due to the uncertainty in the obstacle state $\hat{\mathbf{x}}^{o}_{k+1}$  [^1] For notational brevity, we omit the explicit argument of the CBF $h$ when its dependency is clear from the context.. To account for this uncertainty, we enforce safety to the system using the following probabilistic constraints:
@@ IV-A Safety and Feasibility Analysis
In this section, we identify a key issue in CVaR-BF constraint: the tuning hyper-parameter, the risk level $\beta$, presents a trade-off between the safety and feasibility of the optimization problem in (14). Specifically, this risk level $\beta$, which is set by the user and held constant throughout the trajectory [5], plays a critical role in determining the robot’s behavior near the boundary of the safe set. We analyze this issue from a set-based perspective in two aspects: 1) safety and 2) feasibility.
@@ IV-B Adaptive Risk Level
V Dynamic Safety Zones for CVaR Barrier Functions
@@ V-A Dynamic Zone–Based Barrier Function
$h^{\text{D}}_{k}=\|\mathbf{p}_{k}-\mathbf{p}^{o}_{k}\|^{2}-R_{\text{safe}}^{2},$ (21)
where $R_{\text{safe}}$ represents a threshold distance capturing the minimal allowable separation. This formulation may lead to myopic behavior, causing the robot to navigate too close to obstacles and thereby increasing the risk of collision.
@@
$\mathbf{p}^{\text{rel}}_{k}=\mathbf{p}_{k}-\mathbf{p}^{o}_{k},\quad\mathbf{v}^{\text{rel}}_{k}=\mathbf{v}_{k}-\mathbf{v}^{o}_{k},$
where $\phi$ is the half-angle of the collision cone, defined as $\cos\phi=\sqrt{\|\mathbf{p}^{\text{rel}}_{k}\|^{2}-R_{\text{safe}}^{2}}\,/\,\|\mathbf{p}^{\text{rel}}_{k}\|$. This CBF enforces that the angle between $\mathbf{p}^{\text{rel}}_{k}$ and $\mathbf{v}^{\text{rel}}_{k}$ remains less than $\pi-\phi$, thereby ensuring that the robot is directed away from the obstacle. However, such geometric constraints can be overly conservative. In highly dynamic environments, they may force the robot to execute unnecessarily, leading to premature path diversions and, in some cases, rendering the navigation problem infeasible [12].
@@
$h^{\text{Z}}_{k}$ $\coloneqq\|\mathbf{p}_{k}-\mathbf{p}^{o}_{k}\|^{2}-R_{\text{safe}}^{2}\Bigl{(}1+\Delta_{k}\Bigr{)},$ (23)
@@ V-B Probabilistic Safety Guarantee
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}\geq(1-\gamma)\,h^{\text{Z}}_{k}\Rightarrow$ (27)
@@
$\operatorname{CVaR}^{k}_{\bar{\beta}_{u}}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}=$ $\operatorname{CVaR}^{k}_{\beta_{u}}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}+$ (28)
@@ V-C Risk Adaptive CVaR-BF Optimization
To approximate the CVaR term in the safety constraint, we sample noise $\mathbf{w}_{k}$ in a finite set $\mathcal{W}=\{\mathbf{w}^{(1)},\mathbf{w}^{(2)},\dots,\mathbf{w}^{(L)}\}$, where each $\mathbf{w}^{(j)}\in\mathbb{R}^{n}$ occurs with probability $p\bigl{(}\mathbf{w}^{(j)}\bigr{)}\coloneqq P\Bigl{(}\mathbf{w}_{k}=\mathbf{w}^{(j)}\Bigr{)},\,j=1,2,\dots,L$. Assume $\mathbf{w}_{k}$ is independent of the deterministic initial condition for all $k$. For each noise sample, we compute the corresponding value of the random variable $h_{k+1}$, which represents the safety measure at time $k+1$. These samples are then used to approximate the expectation in the CVaR formulation in (9). Using the CVaR definition from (9) and its reformulation in (14), the following optimization problem is solved at each time $k$:
@@
$\quad\quad-\Bigl{(}\zeta_{i}+\frac{1}{\beta_{k}}\sum_{j=1}^{L}p_{j}\,\eta_{j}\Bigr{)}\geq(1-\gamma_{i})\,-h^{\text{Z}}_{i,{k}},$
@@ VI-A2 Agent Settings
We let the obstacles follow the Social Force Model (SFM) [28] with uncooperative behavior, meaning it only avoids collisions with other obstacles. Their maximum speeds along each axis are chosen uniformly from $\{0.3,0.6,0.9,1.2\}$m/s, and their radii are selected from $\{0.3,0.4,0.5\}$m. The robot is modeled as a double integrator, with dynamics given by
@@ VI-B Benchmark Comparisons
We provide more analysis of the statistical SR in Fig. 4 with respect to different number of obstacles and noise levels. Fig. 4a shows that in deterministic environment ($\sigma=0.0$), CrowdNav++ achieves the highest success rate since it is trained in the same deterministic environment. CrowdNav lacks obstacle trajectory prediction, thus exhibiting a lower success rate. Note that all CVaR-BF based methods reduce to standard CBF methods when $\sigma=0.0$, leading to ineffective risk adjustments, therefore does not show superior performance. However, since our approach incorporates a dynamic zone, it still outperforms cone-based and distance-based barrier functions. As shown in Fig. 4b, introducing noise reduces the SR of RL methods regardless of the number of obstacles. Although CrowdNav++ considers randomized obstacle behaviors (e.g., goal change and size variation), it still experiences a significant performance drop under uncertainty. More conservative methods, e.g., RCBF(Cone) and CVaR(Dist) with $\beta=0.01$ also suffer significant degradation due to feasibility issues, resulting in the worst performance. An interesting finding is that our method performs even better in uncertain environments (Fig. 4b) than in deterministic ones (Fig. 4a). This is because in deterministic environments, our method degrades to CBF, thus the entire risk adaptation mechanism becomes void, making it less flexible. Fig. 4b and Fig. 4c further indicate that as noise level increases, the performance of all methods deteriorates. However, our proposed approach consistently achieves the highest success rate, demonstrating its robustness against uncertainty. More results can be found in the video here.

### v3

@@ Abstract
Robot navigation in dynamic, crowded environments poses a significant challenge due to the inherent uncertainties in the obstacle model. In this work, we propose a risk-adaptive approach based on the Conditional Value-at-Risk Barrier Function (CVaR-BF), where the risk level is automatically adjusted to accept the minimum necessary risk, achieving a good performance in terms of safety and optimization feasibility under uncertainty. Additionally, we introduce a dynamic zone-based barrier function which characterizes the collision likelihood by evaluating the relative state between the robot and the obstacle. By integrating risk adaptation with this new function, our approach adaptively expands the safety margin, enabling the robot to proactively avoid obstacles in highly dynamic environments. Comparisons and ablation studies demonstrate that our method outperforms existing social navigation approaches, and validate the effectiveness of our proposed framework. [Paper Page][Video][Code].
@@ III-A Discrete-Time Control Barrier Functions
where $\mathbf{x}_{k}\in\mathcal{X}\subset\mathbb{R}^{n}$ is the state at time step $k\in\mathbb{Z}^{+}$, and $\mathbf{u}_{k}\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input, with $\mathcal{U}$ being the set of admissible controls for system (1). The function $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is assumed to be locally Lipschitz continuous with respect to $\mathbf{x}_{k}$ and $\mathbf{u}_{k}$.
@@ III-B Probabilistic Constraints
Due to the inherent uncertainty in the dynamic obstacle scenarios, the safe states of the robot cannot be computed deterministically. Therefore, a probabilistic formulation of safety is required. The estimated obstacle state at time $k$, denoted $\hat{\mathbf{x}}^{o}_{k}\in\mathcal{O}\subset\mathbb{R}^{n}$, is obtained via sensor data and state estimation, and is treated as a deterministic value in our framework. However, $\mathbf{x}^{o}_{k+1}$ is modeled as a random variable, reflecting the inherent uncertainty in the predicted motion of the obstacle. Define the safe set as
$\mathcal{S}=\{\mathbf{x}_{k}\in\mathcal{X}:h(\mathbf{x}_{k},{\color[rgb]{0,0,0}\hat{\mathbf{x}}^{o}_{k})}\geq 0\}.$ (5)
$h:\mathcal{X}\times\mathcal{O}\to\mathbb{R}$ is the CBF that depends on both the system state $\mathbf{x}_{k}$ and the estimated obstacle state $\hat{\mathbf{x}}^{o}_{k}$ at the current time step $k$. Therefore, the CBF at the next time step $k+1$ considering the predicted state, which is denoted as $h_{k+1}\coloneqq h\big{(}\mathbf{x}_{k+1},\mathbf{x}^{o}_{k+1}\big{)},$ becomes a random variable due to the uncertainty in the obstacle state $\mathbf{x}^{o}_{k+1}$  [^1] For notational brevity, we omit the explicit argument of the CBF $h$ when its dependency is clear from the context..
To account for this uncertainty, we enforce safety to the system using the following probabilistic constraints:
@@ IV-A Safety and Feasibility Analysis
We identify a key issue in CVaR-BF constraint: the tuning hyper-parameter, the risk level $\beta$, presents a trade-off between the safety and feasibility of the optimization problem in (13). Specifically, this risk level $\beta$, which is set by the user and held constant throughout the trajectory [5], plays a critical role in determining the robot’s behavior near the boundary of the safe set. We analyze this issue from a set-based perspective in two aspects: safety and feasibility.
@@ IV-B Adaptive Risk Level
V Dynamic Zones for CVaR Barrier Functions
@@ V-A Dynamic Zone–Based Barrier Function
$h^{\text{D}}_{k}=\|\mathbf{p}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}}\|^{2}-R_{\text{safe}}^{2},$ (20)
where $R_{\text{safe}}$ represents a threshold distance capturing the minimal allowable separation, and $\mathbf{p}_{k}\in\mathbb{R}^{d}$ and ${\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}}\in\mathbb{R}^{d}$ represent the position of robot and obstacle w.r.t. a global frame (with $n=2d$ for a $d$-dimensional space), respectively. This formulation may lead to myopic behavior, causing the robot to navigate too close to obstacles and thereby increasing the risk of collision.
@@
$\mathbf{p}^{\text{rel}}_{k}=\mathbf{p}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}},\quad\mathbf{v}^{\text{rel}}_{k}=\mathbf{v}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{v}}^{o}_{k}},$
where $\phi$ is the half-angle of the collision cone, defined as $\cos\phi=\sqrt{\|\mathbf{p}^{\text{rel}}_{k}\|^{2}-R_{\text{safe}}^{2}}\,/\,\|\mathbf{p}^{\text{rel}}_{k}\|$, and $\mathbf{v}_{k}\in\mathbb{R}^{d}$ and ${\color[rgb]{0,0,0}\hat{\mathbf{v}}^{o}_{k}}\in\mathbb{R}^{d}$ represent the velocity of robot and obstacle, respectively. This CBF enforces that the angle between $\mathbf{p}^{\text{rel}}_{k}$ and $\mathbf{v}^{\text{rel}}_{k}$ remains less than $\pi-\phi$, thereby ensuring that the robot is directed away from the obstacle. However, such geometric constraints can be overly conservative. In highly dynamic environments, they may force the robot to execute unnecessarily, leading to premature path diversions and, in some cases, rendering the navigation problem infeasible [12].
@@
$h^{\text{Z}}_{k}$ $\coloneqq\|\mathbf{p}_{k}-{\color[rgb]{0,0,0}\hat{\mathbf{p}}^{o}_{k}}\|^{2}-R_{\text{safe}}^{2}\Bigl{(}1+\Delta_{k}\Bigr{)},$ (22)
@@ V-B Probabilistic Safety Guarantee
$\operatorname{CVaR}^{k}_{\beta}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}\geq(1-\gamma)\,h^{\text{Z}}_{k}\Leftrightarrow$ (26)
@@
$\operatorname{CVaR}^{k}_{\bar{\beta}_{u}}\bigl{(}h^{\text{Z}}_{k+1}\bigr{)}=$ $\operatorname{CVaR}^{k}_{\beta_{u}}\bigl{(}h^{\text{D}}_{k+1}\bigr{)}+$ (27)
@@ V-C Risk Adaptive CVaR-BF Optimization
To approximate the CVaR term in the safety constraint, for each obstacle, we consider a finite set of possible realizations of the next obstacle state, $\mathcal{X}^{o}_{k+1}=\{\mathbf{x}^{o,(1)}_{k+1},\mathbf{x}^{o,(2)}_{k+1},\dots,\mathbf{x}^{o,(L)}_{k+1}\}$, each associated with a probability $p(\mathbf{x}^{o,(j)}_{k+1})$. These samples may be drawn from any predictive distribution or model that captures the uncertainty in the obstacle’s next state. For each realization, we compute the corresponding value of the random variable $h_{k+1}$, which represents the safety measure at time $k+1$. These samples are then used to approximate the expectation in the CVaR formulation in (8). Using the CVaR definition from (8) and its reformulation in (13), the following optimization problem is solved at each time $k$:
@@
$\quad\quad-\Bigl{(}\zeta_{i}+\frac{1}{\beta_{k}}\sum_{j=1}^{L}p_{j}\,\eta_{j}\Bigr{)}\geq(1-\gamma_{i})\,h^{\text{Z}}_{i,{k}},$
@@
Remark 2.
Feasibility condition: By setting $\beta=1$, Problem 1 reduces to the deterministic CBF-based optimization with multiple constraints and input limits. Therefore, as long as the deterministic problem is feasible, Problem 1 is always guaranteed to be feasible by setting $\beta=1$, even if no feasible solution exists for any $\beta<1$.
@@ VI-A2 Agent Settings
We assess our method in a widely used crowd navigation simulator within a 12m $\times$ 12m space [28] (see Fig. 3), where each obstacle follows the Social Force Model (SFM) [29] with uncooperative behavior, meaning obstacles avoid collisions only with one another, not with the robot. Their maximum speeds along each axis are chosen uniformly from $\{0.3,0.6,0.9,1.2\}$m/s, and their radii are selected from $\{0.3,0.4,0.5\}$m. The robot is modeled as a double integrator, with dynamics given by
@@ VI-B Benchmark Comparisons
We provide more analysis of the statistical SR in Fig. 4 with respect to different number of obstacles and noise levels. Fig. 4a shows that in deterministic environment ($\sigma=0.0$), CrowdNav++ achieves the highest success rate since it is trained in the same deterministic environment. CrowdNav lacks obstacle trajectory prediction, thus exhibiting a lower success rate. Note that all CVaR-BF based methods reduce to standard CBF methods when $\sigma=0.0$, leading to ineffective risk adjustments, therefore does not show superior performance. However, since our approach incorporates a dynamic zone, it still outperforms cone-based and distance-based barrier functions. As shown in Fig. 4b, introducing noise reduces the SR of RL methods regardless of the number of obstacles. Although CrowdNav++ considers randomized obstacle behaviors (e.g., goal change and size variation), it still experiences a significant performance drop under uncertainty. More conservative methods, e.g., RCBF(Cone) and CVaR(Dist) with $\beta=0.01$ also suffer significant degradation due to feasibility issues, resulting in the worst performance. An interesting finding is that our method performs even better in uncertain environments (Fig. 4b) than in deterministic ones (Fig. 4a). This is because in deterministic environments, our method degrades to CBF, thus the entire risk adaptation mechanism becomes void, making it less flexible. Fig. 4b and Fig. 4c further indicate that as noise level increases, the performance of all methods deteriorates. However, our proposed approach consistently achieves the highest success rate, demonstrating its robustness against uncertainty.
VI-C Ablation Studies